OpenStudy (anonymous):

how many significant digit of 0.03 note:it is about relative error Note that the tail portion of the form 5000 . . . = 4999 . . . is still allowed .EXAMPLES 5.1 has 1 significant digit of 5: |5 − 5.1| = 0.1 0.51 has 1, not 2, significant digits of 0.5: |0.5 − 0.51| = 0.01 4.995 has 3 significant digits of 5: 5 − 4.995 = 0.005 4.994 has 2, not 3, significant digits of 5: 5 − 4.994 = 0.006 0.5 has all significant digits of 0.5 1.4 has 0 significant digits of 2: 2 − 1.4 = 0.6

OpenStudy (anonymous):

This may go better in the Chemistry or Physics section in the future should you have questions like this. But the answer is 1.

OpenStudy (anonymous):

1

OpenStudy (unklerhaukus):

zeros to the left of the first nonzero number are insignificant

OpenStudy (anonymous):

it is about relative error Note that the tail portion of the form 5000 . . . = 4999 . . . is still allowed .EXAMPLES 5.1 has 1 significant digit of 5: |5 − 5.1| = 0.1 0.51 has 1, not 2, significant digits of 0.5: |0.5 − 0.51| = 0.01 4.995 has 3 significant digits of 5: 5 − 4.995 = 0.005 4.994 has 2, not 3, significant digits of 5: 5 − 4.994 = 0.006 0.5 has all significant digits of 0.5 1.4 has 0 significant digits of 2: 2 − 1.4 = 0.6

OpenStudy (anonymous):

my teacher said 3 digit

OpenStudy (unklerhaukus):

i dont think that is right at all

OpenStudy (noelgreco):

Ask your teacher this: Millikan found that the charge on an electron was .00000000000000000016 Coulomb. Did he measure this to the nearer ten quintillionth?

OpenStudy (nincompoop):